OpenStudy (anonymous):

y = x2 + 4x − 21 and how to find both. Keep in mind, your goal is to help Patricia understand the "concepts", not just how to use the steps.

OpenStudy (mimi_x3):

(x-3)(x+7) x = 3 , x = -7

OpenStudy (saifoo.khan):

Someone cant read the english.

OpenStudy (saifoo.khan):

the question says, HELP HER UNDERSTAND THE CONCEPT.

OpenStudy (mimi_x3):

I can read the english , I KNOW RELAX , but I suck at explaining

OpenStudy (mimi_x3):

well let me try a+b = -3 + 7 a*b = -3*7 So you substitute it in , so its (x-3)(x+7)

OpenStudy (saifoo.khan):

Lol, i was jk.

OpenStudy (anonymous):

lol and thanks

OpenStudy (mimi_x3):

lols , i thought yu were serious xD gosh , im gullible

OpenStudy (saifoo.khan):

Srsly u r!

OpenStudy (mimi_x3):

me ? Im not , I was xD

OpenStudy (saifoo.khan):

Haha, NOOO!

OpenStudy (mimi_x3):

NOO ! I shoud know , I'm not gullible

OpenStudy (saifoo.khan):

NOO ! I shoud know , You r gullible

OpenStudy (mimi_x3):

plagiarism detected again

OpenStudy (saifoo.khan):

who, u got a built-in detector!

OpenStudy (mimi_x3):

YUP ! xD

OpenStudy (saifoo.khan):

wow, how much that cost?

OpenStudy (mimi_x3):

we should stop incase the person gets annoyed again like last time

OpenStudy (saifoo.khan):

i will buy two!

OpenStudy (saifoo.khan):

LOL. i love to annoy ppl.

OpenStudy (mimi_x3):

hm , its priceless

OpenStudy (saifoo.khan):

(till they get fulli angry)

OpenStudy (mimi_x3):

lols , annoy people like me ? xD

OpenStudy (saifoo.khan):

fully*

OpenStudy (mimi_x3):

lmao , and delete the question xD

OpenStudy (saifoo.khan):

Wow, i will get 3 then!

OpenStudy (saifoo.khan):

HAHA.

OpenStudy (mimi_x3):

its priceless , you cant buy it anywhere hehe

OpenStudy (mimi_x3):

only i have possession to it xD

OpenStudy (saifoo.khan):

OMg! please give that to me.. i will pay u the rent.

OpenStudy (mimi_x3):

lmao , you can't afford rent :P

OpenStudy (saifoo.khan):

How much is it? i can get that for some minutes though

OpenStudy (mimi_x3):

lmao , its priceless , you cant even get hold to it hehe

OpenStudy (saifoo.khan):

Can i even see that device??

OpenStudy (mimi_x3):

nahh :P It's too valuable to be seen